Grade 12Algebra

for the quadratic equation,ax2+bx+c=0,find the condition such that one root is nth power of the other root.

Nishtha Gahlot
9 Years agoGrade 12

1 Answer

Saurabh Koranglekar
6 Years ago

To determine the condition for a quadratic equation of the form ax² + bx + c = 0 where one root is the nth power of the other root, we can start by letting the roots be represented as r and r^n. Here, r is one root, and r^n is the other root. We can use Vieta's formulas, which relate the coefficients of the polynomial to sums and products of its roots, to derive the necessary condition.

Using Vieta's Formulas

According to Vieta's formulas for a quadratic equation, if r and r^n are the roots, then:

  • The sum of the roots (r + r^n) is equal to -b/a.
  • The product of the roots (r * r^n) is equal to c/a.

Setting Up the Equations

Based on Vieta's, we can express the following two equations:

  1. r + r^n = -b/a
  2. r * r^n = c/a

Rearranging the Product of Roots

From the second equation, we can rewrite it as:

r^(n + 1) = c/a.

This tells us that the product of the roots is dependent on the coefficient c and the leading coefficient a.

Substituting r in the Sum of Roots

Next, we can manipulate the first equation. Let's express r in terms of c/a. From the product equation, we find:

r = (c/a)^(1/(n + 1)).

Now substituting r back into the sum equation:

(c/a)^(1/(n + 1)) + (c/a)^(n/(n + 1)) = -b/a.

Finding the Common Basis

Notice that both terms on the left can be expressed with a common base:

(c/a)^(1/(n + 1)) * (1 + (c/a)^(n/(n + 1) - 1)) = -b/a.

Deriving the Condition

The expression can be simplified further, but it essentially leads us to a relationship involving the coefficients a, b, and c, specifically in terms of the parameter n. For practical purposes, the key insight here is that:

  • The values of a, b, and c must be such that this equation holds true.
  • This often results in a specific ratio or relationship among these coefficients depending on the value of n.

Example for Clarity

Let's take a specific case where n = 2. If we simplify the conditions:

  • We have r + r² = -b/a.
  • Also, r * r² = r³ = c/a.

From this, we can derive a specific condition for values of a, b, and c which satisfy this relationship. You would substitute values or find a particular relationship (like a quadratic in terms of r) that holds for these roots.

In summary, the condition for one root being the nth power of the other in a quadratic equation can be derived from Vieta's formulas, leading to a specific relationship involving the coefficients of the equation. The exact nature of this relationship will depend on the value of n and can be expressed in several forms based on the requirements of the problem at hand.